Find the values of x and y
4x-5y=-3
9x-2y=-6

Answer by user_dude2008(1861)

4x-5y=-3
9x-2y=-6


2*(4x-5y)=-3*2
-5(9x-2y)=-6*-5


8x-10y=-6
-45x+10y=30
-----------
-37x=24
x=-24/37


4x-5y=-3
4(-24/37)-5y=-3
-96-185y=-111
-185y=-15
y=3/37

Answers: x=-24/37, y=3/37